OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.) for this one i havew 1/2pi - 6 inches^2

OpenStudy (anonymous):

from one corner of the square to the opposite corner is a diameter of the circle so using pythagoras you have d=6sqrt(2) so the radius is 3sqrt(2) now you can find the area of both the square and the circle the square being 36 in^2 and the circle being 6pi so one segment will be the sum of the difference of the area between the cirlce and the square all divided by 4 so (72pi-36)/4=36(2pi-1)/4=9(2pi-1)